# Orbit-based structural decomposition and stellar population recovery for edge-on barred galaxies

### v1

@@ 5.4 Stellar metallicity profiles and mean stellar metallicities of different structures
Similar to Fig. 12, the mass-weighted stellar metallicity profiles for different structures are shown in Fig. 14. For the bars and bulges, our models successfully reproduce their steep negative metallicity gradients and can distinguish their metallicity differences. For the disks and halos, our models generally recover their flat or weakly positive metallicity gradients within the inner regions ($R\lesssim 30\,\rm arcsec$; $\rm 1\,arcsec=0.2\,kpc$) and their flat or weakly negative gradients in the outer regions ($R\gtrsim 30\,\rm arcsec$).
Similar to Fig. 13, we compare the true and model-predicted mass-weighted mean stellar metallicities within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume in Fig. 15. The mean metallicity biases $Z_{\rm model}-Z_{\rm true}$ are 0.1, 0.3, 0.1, and $\rm 0.1\,Z_{\odot}$, respectively. The models maintain the absolute biases of the stellar metallicities within $\rm\sim 0.5\,Z_{\odot}$ for the bar, disk, and halo components. However, the bulge metallicities of Au-18 tend to be overestimated (up to $1.1\,Z_{\odot}$). This discrepancy is mainly caused by the models’ underestimation of the bulge mass in outer regions ($R\gtrsim 10\,\rm arcsec$), where the stellar populations are significantly more metal-poor than those in the central bulge regions ($R\lesssim 10\,\rm arcsec$).
@@ 6.1 Method improvements and remaining challenges
In this study, we combine the barred orbit-superposition method (Tahmasebzadeh et al., 2021, 2022) with the population-orbit superposition method (Zhu et al., 2020), and extend the combined new method to model edge-on barred galaxies with effective improvements. By incorporating both the morphologies and kinematics of stellar orbits into the structural decomposition framework, our method achieves higher accuracy in recovering different structures ($|f_{\rm model}-f_{\rm true}|\lesssim 0.1$) compared to earlier non-edge-on barred galaxy models (Tahmasebzadeh et al., 2022), where the biases for bar and bulge fractions reached $\sim 0.15$. Additionally, when tagging the stellar orbits with ages and metallicities, we do not use the age-circularity and age-metallicity correlations adopted in previous works (Zhu et al., 2020; Ding et al., 2023; Jin et al., 2024), which avoid strong assumptions in the modelling. Despite weaker model constraints, our method still provides robust recoveries for stellar populations in different components, which reduces the reliance on prior assumptions and allows flexible studies.
As shown in Fig. 11, the mass fractions of bar and disk exhibit larger biases and uncertainties than the bulge and halo components. This is attributed to uncertainties of the model-predicted dynamical bar lengths $R_{\rm bar,model}$. The values of the true and model-predicted bar lengths are shown in Table 4. For edge-on observations, it is hard to distinguish between the bar end and the disk. Both the bar end and the disk show similar luminosity distributions from edge-on views and are rotation-dominated ($\lambda_{z}\sim 1$; see Fig. 4). The key difference lies in their stellar orbits: disk orbits tend to be round in the $x$-$y$ plane ($p_{\rm orb}\sim 1$), while bar orbits elongate along the bar ($p_{\rm orb}\ll 1$). However, such a difference in orbit morphologies is indistinguishable with data from edge-on views, which results in the uncertainties of dynamical bar lengths.
Previous studies have shown that the presence of a bar produces a high-velocity tail in line-of-sight velocity distributions, corresponding to a positive $h_{3}$$V$ correlation (e.g. Bureau & Athanassoula, 2005; Molaeinezhad et al., 2016; Li et al., 2018), while disk regions exhibit a negative $h_{3}$$V$ correlation. Therefore, the transition between positive and negative $h_{3}$$V$ correlations could indicate the bar length. For example, the true dynamical bar length of Au-23 ($24.6\,\rm arcsec$; see Table 4) is similar to the transition radius estimated from the $h_{3}$ map of Au-23-85-50 ($R_{\rm bar,h3}\sim 22\,\rm arcsec$; see Fig. 1).
@@
6.2 The coexistence of bars, classical bulges, and nuclear disks
@@
The central structures of barred galaxies can be formed by various physical processes. Studies showed that during the secular evolution of galaxies, bars can channel gas towards galaxies’ centres, gradually forming disk-like bulges (e.g. Athanassoula, 2003; Kormendy & Kennicutt, 2004; Gadotti, 2011). Additionally, vertical instabilities in bars will result in BP/X-shaped structures (bar buckling; e.g. Raha et al., 1991; Kuijken & Merrifield, 1995; Athanassoula, 2005; Debattista et al., 2006; Martinez-Valpuesta et al., 2006). Both types are collectively termed ‘pseudo-bulges’, which are dominated by rotation. These contrast with the random-motion-dominated ‘classical bulges’, which are elliptical-like and are likely formed by dissipationless collapse or mergers (e.g. Bournaud et al., 2007; Fisher & Drory, 2008; Gadotti, 2009; Hopkins et al., 2010). Due to different formation mechanisms, classical bulges are typically older and more metal-poor than pseudo-bulges. Whether barred galaxies can host classical bulges has remained controversial for decades.
@@
In our own Galaxy, recent observations have revealed dynamically hot, old, and metal-poor stars in the bulge region. Some of these stars likely originated from the ancient thick disk/halo, while others may belong to a small classical bulge (e.g. Ness et al., 2013; Di Matteo et al., 2015; Rojas-Arriagada et al., 2017; Zoccali et al., 2017; Lucey et al., 2021). This supports the opinion that classical bulges can coexist with bars in galaxies. For external galaxies, some attempts have been made to model classical bulges in barred galaxies (e.g. Méndez-Abreu et al., 2014; Erwin et al., 2015), but based on photometric decompositions that exhibit large degeneracies. For the simulated galaxies analysed in this paper, the bulges identified through stellar orbits are dynamically hotter, older, and more metal-poor than the bars, which are consistent with the properties of classical bulges. Our models successfully reproduce these properties, demonstrating the method’s ability to study the coexistence of bars and classical bulges in external galaxies together with their chemical properties.
@@
Observations show that nuclear disks are common in barred galaxies (e.g. Gadotti et al., 2020; Fraser-McKelvie et al., 2025). Using the barred orbit-superposition method developed by Tahmasebzadeh et al. (2021, 2022), the nuclear disk in a non-edge-on galaxy NGC 4371 was successfully detected (Tahmasebzadeh et al., 2024). Although the simulated galaxies we used in this paper lack nuclear disks, our method can effectively identify such structures if present. Nuclear disks are rotation-dominated ($\lambda_{z}\sim 1$), and their time-averaged orbital radii $R$ are smaller than the bar lengths, which means that these structures are identifiable in the $\lambda_{z}$$R$ phase space.
@@
Overall, our decomposition method allows us to investigate the coexistence of bars, classical bulges, and nuclear disks in real observations. By further analysing the ages and metallicities of these structures, we can better understand their formations.
@@ 7 Summary
We select three barred galaxies from the Auriga simulations and create 12 edge-on mock observations in total (four different projections for each galaxy). The orbit-superposition models for these galaxies have already been constructed in our previous paper (Jin et al., 2025). In this paper, we further decompose these galaxies into bars, bulges, disks, and halos based on stellar orbits. We then tag the stellar orbits with ages and metallicities to derive the stellar populations of each structure. We compare the properties of model-predicted components with their true counterparts. Our main findings are:
@@
    Our models well recover the morphologies, stellar kinematics, ages, and metallicities of different components, including (BP/X-shaped) bars, spheroidal bulges, thin disks, and spatially diffuse stellar halos.
@@
    The mass fractions of all components can be constrained with absolute biases $|f_{\rm model}-f_{\rm true}|\leq 0.15$. For the disk and bar components, the absolute biases reach maximums of 0.15 and 0.14, respectively. For the halo components, the mass fractions are recovered with $|f_{\rm model}-f_{\rm true}|\leq 0.03$. For the bulge components, 10 of 12 cases have $|f_{\rm model}-f_{\rm true}|\leq 0.05$; the other two cases have $|f_{\rm model}-f_{\rm true}|\leq 0.10$.
@@
    Our models generally recover the negative age gradients in the bars and disks of the three simulated galaxies, while the model-predicted age gradients for the bulges and halos have large uncertainties. Nevertheless, the mean stellar ages of all components are constrained with absolute biases $|t_{\rm model}-t_{\rm true}|\rm\lesssim 1\,Gyr$.
@@
    Our models reproduce the steep negative metallicity gradients in the bars and bulges of the three simulated galaxies, and the flat, weakly negative, or weakly positive gradients in the disks and halos. Except for the bulge in Au-18, the mean stellar metallicities of all other components are well constrained with absolute biases $|Z_{\rm model}-Z_{\rm true}|\rm\leq 0.5\,Z_{\odot}$.
Our method enables the structural analysis for real galaxies in the future, such as the 36 Milky Way-like edge-on galaxies in the MUSE/VLT programme GECKOS (van de Sande et al., 2024). This allows us to investigate the possible coexistence of bars, classical bulges, and nuclear disks, and infer their formation histories through stellar population distributions.
@@ Appendix A Dynamical bar lengths and their influence on model results
We present the values of the true and model-predicted bar lengths in Table 4. As described in Sect. 3.2, these dynamical bar lengths are defined as the smallest radius where the cold orbit fraction $f_{\rm cold}\geq 0.5$. We also present the bar lengths directly estimated from observed kinematic maps by eyes, which correspond to the transition radii between positive and negative $h_{3}$$V$ correlations.
Table 1: Dynamical bar lengths: true values, model predictions, and estimations from observed maps
@@
 [^4] From left to right: (1) Galaxy name; (2) true dynamical bar length $R_{\rm bar,true}$; (3) inclination angle $\theta_{\rm T}$ and bar azimuthal angle $\varphi_{\rm T}$; (4) model-predicted dynamical bar length $R_{\rm bar,true}$; (5) bar length $R_{\rm bar,h3}$ estimated from observed kinematic maps by eyes. We note that $\rm 1\,arcsec=0.2\,kpc$.
To detect the influence of dynamical bar lengths on model results, we use the true dynamical bar lengths $R_{\rm bar,true}$ instead of the model-predicted ones $R_{\rm bar,model}$ to separate the bar/bulge and disk/halo components in the modelling, with other parameters in the decomposition framework unchanged. We present the new model results in Figs. 16 to 21. Compared to the original models, the mass profiles of all components align more closely with the true profiles. The biases and uncertainties of the mass fractions are reduced, particularly for the bars and disks. However, the new models show no significant improvement in reproducing the stellar age and metallicity profiles, primarily because the inner disk and outer bar regions ($R\approx 20\,\rm arcsec$) exhibit similar stellar ages and metallicities.
Figure 16: Similar to Fig. 10, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.
Figure 17: Similar to Fig. 11, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.
Figure 18: Similar to Fig. 12, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.
Figure 19: Similar to Fig. 13, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.
Figure 20: Similar to Fig. 14, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.
Figure 21: Similar to Fig. 15, but for model results with true dynamical bar lengths $R_{\rm bar,true}$.

### v4

@@ 5.4 Stellar metallicity profiles and mean stellar metallicities of different structures
Similarly to Fig. 12, the mass-weighted stellar metallicity profiles for different structures are shown in Fig. 14. For the bars and bulges, our models successfully reproduced their steep negative metallicity gradients and can distinguish their metallicity differences. For the discs and halos, our models generally recovered their flat or weakly positive metallicity gradients within the inner regions ($R\lesssim 30\,\rm arcsec$) and their flat or weakly negative gradients in the outer regions ($R\gtrsim 30\,\rm arcsec$).
We compare the true and model-predicted mass-weighted mean stellar metallicities within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume in Fig. 15, in a similar way to Fig. 13. The mean metallicity biases $Z_{\rm model}-Z_{\rm true}$ are 0.1, 0.3, 0.1, and $\rm 0.1\,Z_{\odot}$, respectively. The models maintain the absolute biases of the stellar metallicities within $\rm\sim 0.5\,Z_{\odot}$ for the bar, disc, and halo components. However, the bulge metallicities of Au-18 tend to be overestimated (up to $1.1\,Z_{\odot}$). This discrepancy is mainly caused by the models’ underestimation of the bulge mass in outer regions ($R\gtrsim 10\,\rm arcsec$), where the stellar populations are significantly more metal-poor than those in the central bulge regions ($R\lesssim 10\,\rm arcsec$).
@@ 6.1 Method improvements and remaining challenges
In this study, we combined the barred orbit-superposition method (Tahmasebzadeh et al., 2021, 2022) with the population-orbit superposition method (Zhu et al., 2020) and extended the combined new method to model edge-on barred galaxies with effective improvements. By incorporating both the morphologies and kinematics of stellar orbits into the structural decomposition framework, our method achieves higher accuracy in recovering different structures ($|f_{\rm model}-f_{\rm true}|\lesssim 0.1$), compared to earlier non-edge-on barred galaxy models (Tahmasebzadeh et al., 2022), where the biases for bar and bulge fractions reached $\sim 0.15$. Additionally, when tagging the stellar orbits with ages and metallicities, we did not use the age-circularity and age-metallicity correlations adopted in previous works (Zhu et al., 2020; Ding et al., 2023; Jin et al., 2024). This allowed us to avoid strong assumptions in the modelling. Despite weaker model constraints, our method still provides robust recoveries for stellar populations in different components, which reduces the reliance on prior assumptions and allows flexible studies.
As shown in Fig. 11, the mass fractions of bar and disc exhibit larger biases and uncertainties than the bulge and halo components. This is attributed to uncertainties of the model-predicted dynamical bar lengths, $R_{\rm bar,model}$. The values of the true and model-predicted bar lengths are shown in Table 4. For edge-on observations, it is hard to distinguish between the bar end and the disc. Both the bar end and the disc show similar luminosity distributions from edge-on views and are rotation-dominated ($\lambda_{z}\sim 1$; see Fig. 4). The key difference lies in their stellar orbits: disc orbits tend to be round in the $x$-$y$ plane ($p_{\rm orb}\sim 1$), while bar orbits elongate along the bar ($p_{\rm orb}\ll 1$). However, such a difference in orbit morphologies is indistinguishable with data from edge-on views, which results in the uncertainties we tend to see on dynamical bar lengths.
Previous studies have shown that the presence of a bar produces a high-velocity tail in line-of-sight velocity distributions, corresponding to a positive $h_{3}$$V$ correlation (e.g. Bureau & Athanassoula, 2005; Molaeinezhad et al., 2016; Li et al., 2018), while disc regions exhibit a negative $h_{3}$$V$ correlation. Therefore, the transition between positive and negative $h_{3}$$V$ correlations could indicate the bar length. For example, the true dynamical bar length of Au-23 ($24.6\,\rm arcsec$; see Table 4) is similar to the transition radius estimated from the $h_{3}$ map of Au-23-85-50 ($R_{\rm bar,h3}\sim 22\,\rm arcsec$; see Fig. 1).
@@
6.2 The coexistence of bars, classical bulges, and nuclear discs
@@
The central structures of barred galaxies can be formed by various physical processes. Studies have shown that during the secular evolution of galaxies, bars can channel gas towards galaxies’ centres, gradually forming disc-like bulges (e.g. Athanassoula, 2003; Kormendy & Kennicutt, 2004; Gadotti, 2011). Additionally, vertical instabilities in bars will result in BP/X-shaped structures (bar buckling; e.g. Raha et al., 1991; Kuijken & Merrifield, 1995; Athanassoula, 2005; Debattista et al., 2006; Martinez-Valpuesta et al., 2006). Both types are collectively termed ‘pseudo-bulges’, which are dominated by rotation. These contrast with the random-motion-dominated ‘classical bulges’, which are elliptical-like and are likely formed by dissipationless collapse or mergers (e.g. Bournaud et al., 2007; Fisher & Drory, 2008; Gadotti, 2009; Hopkins et al., 2010). Due to different formation mechanisms, classical bulges are typically older and more metal-poor than pseudo-bulges. The question of whether barred galaxies can host classical bulges is a decades-long debate.
@@
In our own Galaxy, recent observations have revealed dynamically hot, old, and metal-poor stars in the bulge region. Some of these stars likely originated from the ancient thick disc or halo, while others may belong to a small classical bulge (e.g. Ness et al., 2013; Di Matteo et al., 2015; Rojas-Arriagada et al., 2017; Zoccali et al., 2017; Lucey et al., 2021). This supports the opinion that classical bulges can coexist with bars in galaxies. For external galaxies, some attempts have been made to model classical bulges in barred galaxies (e.g. Méndez-Abreu et al., 2014; Erwin et al., 2015), on the basis, however, of photometric decompositions that tend to exhibit large degeneracies. For the simulated galaxies analysed in this paper, the bulges identified through stellar orbits are dynamically hotter, older, and more metal-poor than the bars, which are consistent with the properties of classical bulges. Our models successfully reproduced these properties, demonstrating the method’s ability to study the coexistence of bars and classical bulges in external galaxies together with their chemical properties.
@@
Observations have shown that nuclear discs are common in barred galaxies (e.g. Gadotti et al., 2020; Fraser-McKelvie et al., 2025). Four of the 12 edge-on galaxies observed in the GECKOS survey exhibit strong kinematic evidence for the presence of nuclear discs. In particular, the complex $h_{3}$$V$ relation is of interest, as it is anti-correlated, correlated, and anti-correlated again in the nuclear disc, bar, and outer disc regions, respectively (Fraser-McKelvie et al., 2025; see also Chung & Bureau, 2004; Bureau & Athanassoula, 2005). This provides an opportunity for dynamical modelling of galaxies incorporating nuclear discs, constrained by MUSE or MUSE-like data. Using the barred orbit-superposition method developed by Tahmasebzadeh et al. (2021, 2022), the nuclear disc in a non-edge-on galaxy NGC 4371 was successfully detected (Tahmasebzadeh et al., 2024). Although the simulated galaxies we used in this paper lack nuclear discs, our method can effectively identify such structures if present. Nuclear discs are rotation-dominated ($\lambda_{z}\sim 1$) and their time-averaged orbital radii, $R,$ are smaller than the bar lengths, which means that these structures are identifiable in the $\lambda_{z}$$R$ phase space.
@@
Overall, our decomposition method allows us to investigate the possible coexistence of bars, classical bulges, and nuclear discs in real observations. By further analysing the ages and metallicities of these structures, we can better understand their formation processes.
@@ 7 Summary
We selected three barred galaxies from the Auriga simulations and created 12 edge-on mock observations in total (four different projections for each galaxy). The orbit-superposition models for these galaxies were constructed in an earlier work and described in our previous paper (Jin et al., 2025). In the current work, we further decomposed these galaxies into bars, bulges, discs, and halos based on stellar orbits. We then tagged the stellar orbits with ages and metallicities to derive the stellar populations of each structure. We compared the properties of model-predicted components with their true counterparts. Our main findings are listed below.
@@
    Our models aptly recovered the morphologies, stellar kinematics, ages, and metallicities of different components, including (BP/X-shaped) bars, spheroidal bulges, thin discs, and spatially diffuse stellar halos.
@@
    The mass fractions of all components can be constrained with absolute biases of $|f_{\rm model}-f_{\rm true}|\leq 0.15$. For the disc and bar components, the absolute biases reach maximums of 0.15 and 0.14, respectively. For the halo components, the mass fractions were recovered with $|f_{\rm model}-f_{\rm true}|\leq 0.03$. For the bulge components, 10 out of 12 cases were characterised by $|f_{\rm model}-f_{\rm true}|\leq 0.05$. The other two cases exhibit $|f_{\rm model}-f_{\rm true}|\leq 0.10$.
@@
    Our models generally recovered the negative age gradients in the bars and discs of the three simulated galaxies, while the model-predicted age gradients for the bulges and halos come with large uncertainties. Nevertheless, the mean stellar ages of all components have been constrained with absolute biases of $|t_{\rm model}-t_{\rm true}|\rm\lesssim 1\,Gyr$.
@@
    Our models reproduced the steep negative metallicity gradients in the bars and bulges of the three simulated galaxies and the flat, weakly negative, or weakly positive gradients in the discs and halos. Apart from the bulge in Au-18, the mean stellar metallicities of all other components are well constrained with absolute biases of $|Z_{\rm model}-Z_{\rm true}|\rm\leq 0.5\,Z_{\odot}$.
Our method will be applicable to structural analyses of real galaxies in the future, such as the 36 Milky Way-like edge-on galaxies in the MUSE/VLT programme GECKOS (van de Sande et al., 2024). These applications will enable investigations of the possible coexistence of bars, classical bulges, and nuclear discs, as well as their formation histories via stellar population distributions.
@@ Appendix A Dynamical bar lengths and their influence on model results
We present the values of the true and model-predicted bar lengths in Table 4. As described in Sect. 3.2, these dynamical bar lengths are defined as the smallest radius where the cold orbit fraction, $f_{\rm cold}\geq 0.5$. We also present the bar lengths directly estimated from observed kinematic maps by eyes, which correspond to the transition radii between positive and negative $h_{3}$$V$ correlations.
Table 1: Dynamical bar lengths: true values, model predictions, and estimations from observed maps [^4] From left to right: (1) galaxy name; (2) true dynamical bar length, $R_{\rm bar,true}$; (3) inclination angle, $\theta_{\rm T}$, and bar azimuthal angle, $\varphi_{\rm T}$; (4) model-predicted dynamical bar length, $R_{\rm bar,model}$; (5) bar length, $R_{\rm bar,h3}$, estimated from observed kinematic maps by eye. We note that $\rm 1\,arcsec=0.2\,kpc$.
@@
To detect the influence of dynamical bar lengths on model results, we use the true dynamical bar lengths, $R_{\rm bar,true}$, instead of the model-predicted ones, $R_{\rm bar,model}$, to separate the bar+bulge and disc+halo components in the modelling, while keeping other parameters in the decomposition framework unchanged. We present the new model results in Figs. 16 to 21. Compared to the original models, the mass profiles of all components align more closely with the true profiles. The biases and uncertainties of the mass fractions are reduced, particularly for the bars and discs. However, the new models show no significant improvement in reproducing the stellar age and metallicity profiles, primarily because the inner disc and outer bar regions ($R\approx 20\,\rm arcsec$) exhibit similar stellar ages and metallicities.
Figure 16: Similarly to Fig. 10, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.
Figure 17: Similarly to Fig. 11, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.
Figure 18: Similarly to Fig. 12, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.
Figure 19: Similarly to Fig. 13, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.
Figure 20: Similarly to Fig. 14, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.
Figure 21: Similarly to Fig. 15, but for model results with true dynamical bar lengths, $R_{\rm bar,true}$.